OpenStudy (anonymous):

Compute the double integral over the region D where D is the triangle with vertices (0,0), (1,1) and (0,1). {D} x\sqrt{y^2-x^2}\, dA

OpenStudy (zarkon):

 \[\int\limits_{0}^{1}\int\limits_{0}^{y}x\sqrt{y^2-x^2}dxdy=\frac{1}{12}\]